# Empirical Best Prediction of Poverty Indicators via Nested Error Regression with High Dimensional Parameters

### v1

@@ 4 Nested error linear regression model with high dimensional parameters
$Y_{ij}=\beta_{0}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{i}+\gamma_{i}+\epsilon_{ij},\quad j=1,\cdots,N_{i};\quad i=1,\cdots,m,$ (10)
where the area effects $\gamma_{i}\overset{\mathrm{iid}}{\sim}N(0,\sigma_{\gamma}^{2})$ and the errors $\epsilon_{ij}\sim N(0,\sigma_{\epsilon i}^{2})$ are independent. To represent the model at the area level, we define the following vectors and matrices for areas $i=1,\cdots,m$:
@@
Under this model, the response vectors $\mathbf{y}_{i}$, $i=1,\cdots,m$, are independent with $\mathbf{y}_{i}\sim N(\bm{\mu}_{i},\mathbf{V}_{i})$, where:
@@
Here, $\mathbf{1}_{N_{i}}$ is an $N_{i}\times 1$ vector of ones, and $\mathbf{I}_{N_{i}}$ is the $N_{i}\times N_{i}$ identity matrix. Following with Molina and Rao (2010), for $i=1,\cdots,m$, we can decompose $\mathbf{y}_{i}$ into sample and out-of-sample elements $\mathbf{y}_{i}=(\mathbf{y}^{\prime}_{is},\mathbf{y}^{\prime}_{ir})^{\prime}$, and the corresponding decomposition on $\mathbf{X}_{i}$, $\bm{\mu}_{i}$ and $\mathbf{V}_{i}$. When the sample size $n_{i}>0$, the distribution of $\mathbf{y}_{ir}$ given the sample data is
$\mathbf{y}_{ir}|\mathbf{y}_{s}\sim N(\bm{\mu}_{ir|s},\mathbf{V}_{ir|s}),$ (11)
@@
$\mathbf{y}_{ir}=\bm{\mu}_{ir|s}+u_{i}\mathbf{1}_{N_{i}-n_{i}}+\bm{e}_{ir},$ (14)
where $u_{i}$ represents new random effects, and $\bm{\epsilon}_{ir}$ denotes the error terms. These components are independent and satisfy the following distributions:
$u_{i}\sim N(0,\sigma_{\gamma}^{2}(1-B_{i})),\quad i=1,\cdots,m,\quad\text{and}\quad\bm{e}_{ir}\sim N(\mathbf{0}_{N_{i}-n_{i}},\sigma_{\epsilon i}^{2}\mathbf{I}_{N_{i}-n_{i}}).$
This reformulation simplifies the simulation process while maintaining the statistical properties of the original model.
As mentioned before, in practice the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, need to be replaced by suitable estimators $\hat{\bm{\phi}}_{i}=(\hat{\beta}_{0},\hat{\bm{\beta}}^{\prime}_{i},\hat{\sigma}_{\gamma}^{2},\hat{\sigma}^{2}_{\epsilon i})$. Then the variables $Y_{ij}$ are generated from the corresponding estimated normal distributions. Given the practical challenges in directly linking the survey sample to the population and under the assumption that the sampling fraction is negligible, for in-sample areas, the elements $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated from
$Y_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\hat{\sigma}_{\gamma}^{2}\mathbf{1}_{n_{i}}^{\prime}\hat{\mathbf{V}}^{-1}_{is}(\mathbf{y}_{is}-\hat{\beta}_{0}\mathbf{1}_{n_{i}}-\mathbf{X}_{is}\hat{\bm{\beta}}_{i})+u_{i}+e_{ij},$
where $u_{i}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}_{\gamma}^{2}(1-\hat{B}_{i}))$, and $e_{ij}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}^{2}_{\epsilon i})$. If an area $i$ is not sampled, then $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated by bootstrap from
@@
where $\gamma_{i}^{*}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}_{\gamma}^{2})$ and $\epsilon^{*}_{ij}\overset{\mathrm{iid}}{\sim}N(0,\hat{\sigma}^{2}_{\epsilon i})$.
@@ 4.1 Estimation of the high dimensional parameters
In this study, we introduce a data-driven method for estimating the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, which effectively leverages data across areas to provide consistent estimators for the area-specific parameters when the tuning parameters are known. Moreover, our approach offers a significant improvement over Lahiri and Salvati (2023) by drastically reducing computational time while maintaining comparable accuracy and performance. These enhancements make our method both more efficient and practical for the analysis of large-scale datasets. The algorithm is detailed in the following steps.
  • Step 1
    Define $r_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-a_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\bm{\beta}_{i})$, where $q_{i}$ is a known scale parameter, ${\mathbf{y}}_{ls}$ is a $n_{l}\times 1$ vector of the response variable and $\mathbf{X}_{ls}$ denotes the matrix of dimension $n_{l}\times p$ individual-level covariates of the sampled units in area $l$. Obtain the area-specific regression coefficients estimates $\{(\hat{a}_{0i},\hat{\bm{\beta}}_{i}),i=1,\cdots,m\}$ via solving the following system of equations:
    $\sum_{l=1}^{m}\mathbf{X}^{\prime}_{ls(p+1)}q_{i}^{1/2}\psi_{i}(r_{l;i})=0,$ (15)
@@
    where $\psi(r)$ is a known monotone non-decreasing function satisfying $\psi(-\infty)<\psi(0)<\psi(\infty)$, and $\tau_{i}\in T=(0,1)$ is a known tuning parameter. Notably, setting $\tau_{i}=0.5$ with the identify function for $\psi$ would lead to the standard least square estimator of the regression coefficient vector. The tuning parameter $\tau_{i}$ introduces small area-specific effects into the estimation of the regression coefficients $\bm{\beta}_{i}$ and the sampling variances $\sigma^{2}_{\epsilon i}$, enhancing the flexibility and adaptability of the model. Following Chambers and Tzavidis (2006), we assume that $\psi(\cdot)$ is the Huber influence function, with a tuning constant $c=1.345$, ensuring robust performance in the presence of outliers.
  • Step 2
    Define $\mathbf{\tilde{r}}_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-\hat{a}_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\hat{\bm{\beta}}_{i})$ and $\tilde{\mathbf{V}}_{l;i}=\sigma^{2}_{\gamma}\bm{1}_{n_{l}}\bm{1}_{n_{l}}^{\prime}+\sigma^{2}_{\epsilon i}\mathbf{I}_{n_{l}}$. Obtain $\{\hat{\sigma}^{2}_{\epsilon i},i=1,\cdots,m\}$ as a solution of the following system of estimation equations:
    $\sum_{l=1}^{m}\Bigg{[}\psi_{i}(\tilde{\mathbf{r}}_{l;i})^{\prime}\;\tilde{\mathbf{V}}_{l;i}^{-1}\;\frac{\partial\mathbf{V}_{i;i}}{\partial\sigma^{2}_{\epsilon i}}\;\tilde{\mathbf{V}}_{l;i}^{-1}\;\psi_{i}(\tilde{\bm{r}}_{l;i})\;-\text{tr}(\tilde{\mathbf{V}}_{l;i}^{-1}\;\frac{\partial\mathbf{V}_{i;i}}{\partial\sigma^{2}_{\epsilon i}})\Bigg{]}=0$ (17)
    where $i=1,\cdots,m$, and $\mathbf{V}_{i;i}=\sigma^{2}_{\gamma}\bm{1}_{n_{i}}\bm{1}_{n_{i}}^{\prime}+\sigma^{2}_{\epsilon i}\mathbf{I}_{n_{i}}$.
  • Step 3
    Define $\bar{\mathbf{r}}^{\star}=\{\bar{r}_{i}^{\star};i\in s\}^{\prime}$, where $\bar{r}_{i}^{\star}=n_{i}^{-1}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$ and $r_{ij}^{\star}=Y_{ij}-\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}$ with $\hat{\bm{\beta}}_{i}$ from Step 1. We estimate $\beta_{0}$ by taking the grand residual mean: $\hat{\beta}_{0}=1/n\sum_{i\in s}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$. Define $v_{i}^{\star}=\sigma^{2}_{\gamma}+\hat{d}_{i}$, where $\hat{d}_{i}=\hat{\sigma}^{2}_{\epsilon i}/n_{i}$ for $i\in s$. Obtain $\hat{\sigma}^{2}_{\gamma}$ as a solution of the following estimating equation:
    $\sum_{i\in s}\bigg{[}\psi^{2}\bigg{(}(v_{i}^{\star})^{-1/2}\bar{r}_{i}^{\star}\bigg{)}(v_{i}^{\star})^{-1}-(v_{i}^{\star})^{-1}w^{\star}\bigg{]}=0$ (18)
@@ 4.2 Definition and estimation of tuning parameter $\tau_{i}$
Let the subscript $U$ denotes a finite population indicators. Define the following finite population parameters:
$Y_{ij;\tau}$ $=$ $\alpha_{0\tau;U}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{\tau;U};$
@@
for $i=1,\cdots,m$; $j=1,\cdots,N_{i}$. Note that $Y_{ij;\tau}$ and $\tau_{ij}$ are unknown even for sampled units since $\alpha_{0\tau;U}$ and $\bm{\beta}_{\tau;U}$ are unknown finite population parameters. We are interested in estimating the following finite population parameter: $\tau_{i}=1/N_{i}\sum_{j=1}^{N_{i}}\tau_{ij}$.
For the sampled areas, we use the same data-driven method as Lahiri and Salvati (2023) to estimate the tuning parameters, $\tau_{i}$ with $i\in s$. Specifically, for a fine grid $\tau\in T$, we obtain a collection of fitted values for the entire sample:
$\hat{Y}_{ij;\tau}$ $=$ $\hat{\alpha}_{0\tau}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{\tau},\quad i\in s;\quad j=1,\cdots,n_{i},$
using the standard quantile or M-quantile methods, where $\hat{\alpha}_{0\tau}$ and $\hat{\bm{\beta}}_{\tau}$ are estimated intercept and regression coefficients, respectively. For each observation $Y_{ij}$ in the sample, we identify the fitted line that minimizes the prediction error, defined as the difference between $Y_{ij}$ and the predicted value from the fitted regression at $\mathbf{x}_{ij}$. Let $\hat{\tau}_{ij}$ denote the value of $\tau$ in the grid corresponding to this best-fitting line. The variability of $\hat{\tau}_{ij}$ reflects unit-level variability. Provided there are sample observations in area $i$, and a non-informative sampling method has been used to obtain them, one can consider an estimate of the area-$i$-specific tuning parameter is the sample average of $\hat{\tau}_{ij}$ for that area: $\bar{\hat{\tau}}_{i}=n_{i}^{-1}\sum_{j\in s_{i}}\hat{\tau}_{ij}$. Lahiri and Salvati (2023) also introduced an empirical linear best (ELB) predictor for $\tau_{i}$ denoted as $\hat{\tau}_{i}^{\text{ELB}}$, which enhances the performance of $\bar{\hat{\tau}}_{i}$ by leveraging data from all areas. For a detailed explanation, we refer readers to their paper.
@@
$\text{E}[\text{logit}(\tau_{i})]=\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta},$
where $i=1,\ldots,m$. The variance components $\Delta_{i}$ are typically assumed to be known; however, in practice, they need to be estimated. In this study, we employ a smoothing technique to estimate $\Delta_{i}$
$\hat{\Delta}_{i}=\frac{1}{n_{i}}\sum_{i\in s}\sum_{j=1}^{n_{i}}\frac{(\hat{\tau}_{ij}-\bar{\hat{\tau}}_{i})^{2}}{n-m_{s}},\quad i\in s,$
where $m_{s}$ is the number of sampled areas. Here, $\mathbf{\bar{Z}}_{i}$ represents the population mean of relevant auxiliary variables, and $\bm{\eta}$ is the unknown vector of coefficients, estimated by $\hat{\eta}$ that minimizes the following objective function, $\mathbf{Q}(\bm{\eta})$:
$\mathbf{Q}(\bm{\eta})=\sum_{i\in s}\frac{1}{\hat{\Delta}_{i}}\left[\bar{\hat{\tau}}_{i}-\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}\right]^{2}.$
@@
$\hat{\tau}_{i}=\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}},\quad i\notin s.$

### v2

@@ 4 Nested error regression model with high dimensional parameters
$Y_{ij}=\beta_{0i}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{i}+\epsilon_{ij},\quad i=1,\cdots,m;\quad j=1,\cdots,N_{i},$ (10)
where $\beta_{0i}=\beta_{0}+\gamma_{i}$. The area effects $\gamma_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\sigma_{\gamma}^{2})$ and the errors $\epsilon_{ij}\sim\mathcal{N}(0,\sigma_{\epsilon i}^{2})$ are independent. To represent the model at the area level, we define the following vectors and matrices for areas $i=1,\cdots,m$:
@@
Under this model, the response vectors $\mathbf{y}_{i}$, $i=1,\cdots,m$, are independent with $\mathbf{y}_{i}\sim\mathcal{N}(\bm{\mu}_{i},\mathbf{V}_{i})$, where:
@@
Here, $\mathbf{1}_{N_{i}}$ is an $N_{i}\times 1$ vector of ones, and $\mathbf{I}_{N_{i}}$ is the $N_{i}\times N_{i}$ identity matrix. Although the area-specific response vectors are conditionally independent under the model, information is shared across areas through the joint estimation of model parameters (see Section 4.1). In particular, the area-specific regression coefficients are obtained by solving estimating equations that involve data from all sampled areas, and the variance component $\sigma^{2}_{\gamma}$ is common across areas. This structure induces borrowing of strength in estimation and prediction, especially for areas with small sample sizes.
The conditional independence assumption refers to independence given the area-specific random effects and covariates. If additional clustering exists within small areas (e.g., primary sampling units nested within areas) and induces further intra-cluster dependence beyond the area effect, such dependence is not explicitly modeled in the present specification of $\mathbf{V}_{i}$. When cluster identifiers are available and such dependence is non-negligible, a multi-level nested-error model incorporating cluster-level random effects may provide a more appropriate specification.
Following Molina and Rao (2010), for $i=1,\cdots,m$, we can decompose $\mathbf{y}_{i}$ into sample and out-of-sample elements $\mathbf{y}_{i}=(\mathbf{y}^{\prime}_{is},\mathbf{y}^{\prime}_{ir})^{\prime}$, and the corresponding decomposition on $\mathbf{X}_{i}$, $\bm{\mu}_{i}$ and $\mathbf{V}_{i}$. When the sample size $n_{i}>0$, the distribution of $\mathbf{y}_{ir}$ given the sample data is
$\mathbf{y}_{ir}|\mathbf{y}_{s}\sim\mathcal{N}(\bm{\mu}_{ir|s},\mathbf{V}_{ir|s}),$ (11)
@@
$\mathbf{y}_{ir}=\bm{\mu}_{ir|s}+u_{i}\mathbf{1}_{N_{i}-n_{i}}+\bm{\epsilon}_{ir},$ (14)
where $u_{i}$ represents new random effects, and $\bm{\epsilon}_{ir}$ denotes a vector of the error terms. These components are independent and satisfy the following distributions:
$u_{i}\sim\mathcal{N}(0,\sigma_{\gamma}^{2}(1-B_{i})),\quad i=1,\cdots,m,\quad\text{and}\quad\bm{\epsilon}_{ir}\sim\mathcal{N}(\mathbf{0}_{N_{i}-n_{i}},\sigma_{\epsilon i}^{2}\mathbf{I}_{N_{i}-n_{i}}).$
This reformulation simplifies the simulation process while maintaining the statistical properties of the original model. As mentioned before, in practice the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, need to be replaced by suitable estimators $\hat{\bm{\phi}}_{i}=(\hat{\beta}_{0},\hat{\bm{\beta}}^{\prime}_{i},\hat{\sigma}_{\gamma}^{2},\hat{\sigma}^{2}_{\epsilon i})$.
In line with Chapter 10 of Morales et al. (2021), when $\alpha=0$ or $\alpha=1$ with a shifted logarithmic transformation applied to the welfare variable $E_{ij}$ (i.e., $Y_{ij}=\log(E_{ij}+c)$ for some constant $c$), explicit expressions of the EBPs for the HCR and PG can be obtained. Specifically, the EBP of the HCR for individual $j$ in area $i$ is given by
$\hat{F}^{EB}_{0ij}=\Phi(\hat{\alpha}_{ij}),\quad\hat{\alpha}_{ij}=\hat{v}_{ij|s}^{-1/2}(T(z)-\hat{\mu}_{ij|s}),$ (15)
where $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution, $\hat{v}_{ij|s}=\hat{\sigma}_{\gamma}^{2}(1-\hat{B}_{i})+\hat{\sigma}_{\epsilon i}$, and
$\hat{\mu}_{ij|s}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\hat{B}_{i}(\bar{y}_{i}-\hat{\beta}_{0}-\mathbf{\bar{x}}_{i}^{\prime}\hat{\bm{\beta}}_{i}),$
with $\bar{y}_{i}=n_{i}^{-1}\sum_{i=1}^{n_{i}}y_{ij}$ and $\bar{\mathbf{x}}_{i}=n_{i}^{-1}\sum_{i=1}^{n_{i}}\mathbf{x}_{ij}$. When the shifted logarithmic transformation is applied to $E_{ij}$, the EBP of the PG for individual $j$ in area $i$ becomes
$\hat{F}^{EB}_{1ij}=\frac{z+c}{z}\Phi(\hat{\alpha}_{ij})-\frac{1}{z}\exp\{\frac{1}{2}\hat{v}_{ij|s}+\hat{\mu}_{ij|s}\}\Phi(\hat{\alpha}_{ij}-\hat{v}_{ij|s}^{1/2}).$ (16)
However, when $\alpha=2$ or other transformation functions are applied to $E_{ij}$, such as the Box–Cox transformation, explicit expressions for the EBPs of the FGT measures are generally not available. In these cases, Monte Carlo approximations provide a practical solution.
Here, the variables $Y_{ij}$ are generated from the corresponding estimated normal distributions. Given the practical challenges in directly linking the survey sample to the population and under the assumption that the sampling fraction is negligible, for in-sample areas, the elements $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated from
$Y_{ij}=\hat{\beta}_{0}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}+\hat{\sigma}_{\gamma}^{2}\mathbf{1}_{n_{i}}^{\prime}\hat{\mathbf{V}}^{-1}_{is}(\mathbf{y}_{is}-\hat{\beta}_{0}\mathbf{1}_{n_{i}}-\mathbf{X}_{is}\hat{\bm{\beta}}_{i})+u_{i}^{*}+\epsilon_{ij}^{*},$
where $u_{i}^{*}\sim\mathcal{N}(0,\hat{\sigma}_{\gamma}^{2}(1-\hat{B}_{i}))$, and $\epsilon_{ij}^{*}\sim\mathcal{N}(0,\hat{\sigma}^{2}_{\epsilon i})$. If an area $i$ is not sampled, then $\{Y_{ij},j=1,\cdots,N_{i}\}$ are generated by bootstrap from
@@
where $\gamma_{i}^{*}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\hat{\sigma}_{\gamma}^{2})$ and $\epsilon^{*}_{ij}\overset{\mathrm{ind}}{\sim}\mathcal{N}(0,\hat{\sigma}^{2}_{\epsilon i})$.
@@ 4.1 Estimation of the high dimensional parameters
In this study, we introduce a data-driven method for estimating the model parameters $\bm{\phi}_{i}=(\beta_{0},\bm{\beta}^{\prime}_{i},\sigma_{\gamma}^{2},\sigma^{2}_{\epsilon i})$ for $i=1,\cdots,m$, which effectively leverages data across areas. Moreover, our approach offers a significant improvement over Lahiri and Salvati (2023) by drastically reducing computational time while maintaining comparable accuracy and performance. These enhancements make our method both more efficient and practical for the analysis of large-scale datasets.
Throughout the remainder of the paper, we use $\{1,\cdots,m_{s}\}$ to denote sampled areas and $\{m_{s}+1,\cdots,m\}$ to index out-of-sample areas, without risk of confusion. The algorithm is detailed in the following steps:
Step 1:
Define $r_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-\beta_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\bm{\beta}_{i}),\;l=1,\cdots,m_{s}$, where $q_{i}$ is a scale factor chosen to stabilize the Huber estimating equations and to reduce the influence of outliers (see, e.g. Bianchi and Salvati (2015)), ${\mathbf{y}}_{ls}$ is a $n_{l}\times 1$ vector of the response variable and $\mathbf{X}_{ls}$ denotes the matrix of dimension $n_{l}\times p$ individual-level covariates of the sampled units in area $l$.
Then, obtain the area-specific regression coefficients estimates $\{(\hat{\beta}_{0i},\hat{\bm{\beta}}_{i}),i=1,\cdots,m\}$ via solving the following system of equations:
$\sum_{l=1}^{m_{s}}\mathbf{X}^{\prime}_{ls(p+1)}q_{i}^{1/2}\psi_{i}(r_{l;i})=0,\;i=1,\cdots,m,$ (17)
@@
where $\psi(r)$ is a known monotone non-decreasing function satisfying $\psi(-\infty)<\psi(0)<\psi(\infty)$, and $\tau_{i}\in T=(0,1)$ is a known tuning parameter. Notably, setting $\tau_{i}=0.5$ with the identify function for $\psi$ would lead to the standard least square estimator of the regression coefficient vector.
The tuning parameter $\tau_{i}$ plays a key role in introducing small area-specific effects into the estimation of both regression coefficients and sampling variances. In contrast to the traditional NER model, which assumes parameter homogeneity across areas, the proposed specification allows $\tau_{i}$ to vary, thereby inducing heterogeneity in $\bm{\beta}_{i}$ and error variances $\sigma^{2}_{\epsilon i}$. This flexibility allows the model to accommodate variations in socio-economic conditions, survey design, and data quality across small areas. Furthermore, when the estimated tuning parameters $\hat{\tau}_{i}$ are similar across areas, the model naturally simplifies to the conventional NER structure. Consequently, the proposed NERHDP framework attains a desirable balance between flexibility and parsimony, enabling it to adapt to heterogeneous data structures while maintaining stability and strong predictive performance.
Following Chambers and Tzavidis (2006), we assume that $\psi(\cdot)$ is the Huber influence function, with a tuning constant $c=1.345$, ensuring robust performance in the presence of outliers.
Step 2:
Define $\tilde{\mathbf{r}}_{l;i}=q_{i}^{-1/2}(\mathbf{y}_{ls}-\hat{\beta}_{0i}\bm{1}_{n_{l}}-\mathbf{X}_{ls}\hat{\bm{\beta}}_{i})$. We estimate the sampling variance $\{\hat{\sigma}^{2}_{\epsilon i},i=1,\cdots,m\}$ by fitting one-way random effects model to $\{\psi_{i}(\tilde{\mathbf{r}}_{l;i}),\;l=1,\cdots,m_{s}\}$, and obtained the REML estimate of the sampling variance. The model fitting is carried out using the function lmer from the R package lme4.
Step 3:
Define $\bar{\mathbf{r}}^{\star}=\{\bar{r}_{i}^{\star};\;i=1,\cdots,m_{s}\}^{\prime}$, where $\bar{r}_{i}^{\star}=n_{i}^{-1}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$ and $r_{ij}^{\star}=Y_{ij}-\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{i}$ with $\hat{\bm{\beta}}_{i}$ from Step 1. We estimate $\beta_{0}$ by taking the grand residual mean: $\hat{\beta}_{0}=1/n\sum_{i=1}^{m_{s}}\sum_{j=1}^{n_{i}}r_{ij}^{\star}$. Define $v_{i}^{\star}=\sigma^{2}_{\gamma}+\hat{d}_{i}$, where $\hat{d}_{i}=\hat{\sigma}^{2}_{\epsilon i}/n_{i}$ for $i=1,\cdots,m_{s}$. Obtain $\hat{\sigma}^{2}_{\gamma}$ as a solution of the following estimating equation:
$\sum_{i=1}^{m_{s}}\bigg[\psi^{2}\bigg((v_{i}^{\star})^{-1/2}\bar{r}_{i}^{\star}\bigg)(v_{i}^{\star})^{-1}-(v_{i}^{\star})^{-1}w^{\star}\bigg]=0,$ (19)
@@
It is worth noting that the proposed procedure shares a common unit-level modeling structure with the approach of Elbers et al. (2003). In both frameworks, poverty indicators are obtained by generating synthetic or predicted unit-level values based on an estimated regression model and associated stochastic components, and then aggregating these values at the small area level.
However, the conceptual foundations differ. The ELL method constructs synthetic populations using a common regression structure and resampled residual components, without conditioning on observed sample values in the sense of empirical best prediction. By contrast, the NERHDP framework derives empirical best predictors under a nested error regression structure and explicitly exploits the conditional distribution of unobserved units given the observed sample. Furthermore, the proposed model allows regression coefficients and sampling variances to vary across areas through area-specific estimating equations.
This distinction clarifies that, while both approaches are unit-level and simulation-based, the proposed methodology extends the conditional EBP paradigm to a heterogeneous high-dimensional parameter setting.
@@ 4.2 Definition and estimation of tuning parameter $\tau_{i}$
Let the subscript $U$ denotes a finite population indicator. Define the following finite population parameters:
$Y_{ij;\tau}$ $=$ $\beta_{0\tau;U}+\mathbf{x}_{ij}^{\prime}\bm{\beta}_{\tau;U};$
@@
for $i=1,\cdots,m$; $j=1,\cdots,N_{i}$. Note that $Y_{ij;\tau}$ and $\tau_{ij}$ are unknown even for sampled units since $\beta_{0\tau;U}$ and $\bm{\beta}_{\tau;U}$ are unknown finite population parameters. We are interested in estimating the following finite population parameter: $\tau_{i}=1/N_{i}\sum_{j=1}^{N_{i}}\tau_{ij}$. The tuning parameter $\tau_{i}$ plays a pivotal role in the NERHDP framework. It governs the degree of local adaptation of the estimating equations and thus controls the extent to which regression coefficients and sampling variances are allowed to vary across areas. When the $\tau_{i}$ values are similar across areas, the model approximates the traditional homogeneous NER structure. Conversely, heterogeneous $\tau_{i}$ values induce area-specific deviations, enabling the model to flexibly accommodate cross-area heterogeneity in regression relationships and dispersion. Therefore, the estimation of $\tau_{i}$ is not merely a technical component, but a central mechanism through which the proposed method balances flexibility and stability.
For the sampled areas, we use the same data-driven method as Lahiri and Salvati (2023) to estimate the tuning parameters for sampled areas, $\tau_{i}$. Specifically, for a fine grid $\tau\in T=(0,1)$, we obtain a collection of fitted values for the entire sample:
$\hat{Y}_{ij;\tau}$ $=$ $\hat{\beta}_{0\tau}+\mathbf{x}_{ij}^{\prime}\hat{\bm{\beta}}_{\tau},\quad i=1,\cdots,m_{s};\quad j=1,\cdots,n_{i},$
using the standard quantile or M-quantile methods, where $\hat{\beta}_{0\tau}$ and $\hat{\bm{\beta}}_{\tau}$ are estimated intercept and regression coefficients, respectively. For each observation $Y_{ij}$ in the sample, we identify the fitted line that minimizes the prediction error, defined as the difference between $Y_{ij}$ and the predicted value from the fitted regression at $\mathbf{x}_{ij}$. Let $\hat{\tau}_{ij}$ denote the value of $\tau$ in the grid corresponding to this best-fitting line. The variability of $\hat{\tau}_{ij}$ reflects unit-level variability. Provided there are sample observations in area $i$, and a non-informative sampling method has been used to obtain them, one can consider an estimate of the area-$i$-specific tuning parameter is the sample average of $\hat{\tau}_{ij}$ for that area: $\bar{\hat{\tau}}_{i}=n_{i}^{-1}\sum_{j\in s_{i}}\hat{\tau}_{ij}$. Lahiri and Salvati (2023) also introduced an empirical linear best (ELB) predictor for $\tau_{i}$ denoted as $\hat{\tau}_{i}^{\text{ELB}}$, which enhances the performance of $\bar{\hat{\tau}}_{i}$ by leveraging data from all areas. For a detailed explanation, we refer readers to their paper.
@@
$\text{E}[\text{logit}(\tau_{i})]=\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta},\quad i=1,\ldots,m.$
We adopt the logit link because $\tau_{i}\in(0,1)$ by definition and the transformation $\operatorname{logit}(\tau_{i})$ maps this constrained parameter to the real line, avoiding boundary issues and stabilizing estimation. Here $\bar{\mathbf{Z}}_{i}$ denotes the finite-population (area-level) mean of the auxiliary variables for small area $i$, i.e., the population average of the covariates over area $i$, and $\bm{\eta}$ is the unknown vector of coefficients.
The choice of $\bar{\mathbf{Z}}_{i}$ is critical in determining the ability of the unmatched model to capture between-area heterogeneity in the tuning parameters $\tau_{i}$. For out-of-sample areas, where $\tau_{i}$ cannot be estimated directly from unit-level data, informative auxiliary variables in $\bar{\mathbf{Z}}_{i}$ that are associated with structural differences in regression coefficients or sampling variances can improve the estimation of $\tau_{i}$ and enhance predictive accuracy. In contrast, if $\bar{\mathbf{Z}}_{i}$ has limited explanatory power, the resulting estimates will tend to behave more synthetically.
The variance components $\Delta_{i}$ are typically assumed to be known; however, in practice, they need to be estimated. In this study, we employ a smoothing technique to estimate $\Delta_{i}$:
$\hat{\Delta}_{i}=\frac{1}{n_{i}}\sum_{i=1}^{m_{s}}\sum_{j=1}^{n_{i}}\frac{(\hat{\tau}_{ij}-\bar{\hat{\tau}}_{i})^{2}}{n-m_{s}},\quad i=1,\cdots,m_{s}.$
And we estimate $\bm{\eta}$ by $\hat{\bm{\eta}}$ that minimizes the following objective function $Q(\bm{\eta})$:
$Q(\bm{\eta})=\sum_{i=1}^{m_{s}}\frac{1}{\hat{\Delta}_{i}}\left[\bar{\hat{\tau}}_{i}-\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\bm{\eta})}}\right]^{2}.$
@@
$\hat{\tau}_{i}=\frac{\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}}{1+\exp{(\bar{\mathbf{Z}}_{i}^{\prime}\hat{\bm{\eta}})}},\quad i=m_{s}+1,\cdots,m.$
@@
In the simulation studies presented in Section 6, the adaptive estimation of $\tau_{i}$ serves as the principal mechanism through which the NERHDP model accommodates heterogeneous data-generating scenarios. In particular, when regression slopes and/or sampling variances vary across areas, the resulting estimates of $\tau_{i}$ capture this cross-area heterogeneity. This adaptive adjustment is expected to improve predictive performance relative to approaches that impose a common (globally homogeneous) structure across areas.